Physics Shift Report
Start Date: 10-10-2017
End Date: 10-23-2017
Start Run: 10016
End Run: 10085
Physics shifter: Tanaz A. Mohayai
Summary of Runs¶
Run Number | Run Type | Absorber | Momentum | Cooling Channel Fields |
10016 | Reference Run | Empty LH2 | 140MeV | Flip Mode - 2T, 2017-02-7 |
10017 - 10018 | Reference Run - Beam Bump Tuning | Empty LH2 | 140MeV | Flip Mode - 2T, 2017-02-7 |
10019 - 10054 | Special Run | Empty LH2 | 140MeV | Flip Mode - 2T, 2017-02-7 |
10055 - 10062 | Special Run - Tracker Timing Scan | Empty LH2 | 140MeV | Flip Mode - 2T, 2017-02-7 |
10063 - 10077 | Special Run | Empty LH2 | 140MeV | Flip Mode - 2T, 2017-02-7 |
10078 - 10085 | Special Run | Empty LH2 | 170MeV | Flip Mode - 3T, 2017-02-8 |
Elog, CDB, and Reconstruction¶
Elog entries are ok.
CDB uploaded ok with the exception of the following runs:
10021: it has null End Date (as reported in the corresponding elog, Run Plan GUI hung up at end of run due to illegal character in end-run comment)
10021: Start comment is wrong and should be 4 mm (not 3 mm)
10022: Start comment is wrong and should be 4 mm (not 3 mm)
10083: not in the CDB but properly explained in the corresponding MICE log (an attempt to start a Reference run was made but was never properly recorded in CDB)
Online reconstruction is ok
All runs are reconstructing ok with the exception of the following:
10083: was never recorded in CDB (please see above for explanation)
Timing Study¶
There were tracker timing studies carried out during this physics shift period. The corresponding run numbers are in Summary of Runs table above.
It was concluded that the baseline time (the current time) setting for the H40a task in the current run plan (3-140+M3-Test3 beamline tag in https://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20171015) is the optimum. This is because the "the maximum of the accepted trigger fraction is found at this baseline time". See the plot attached to elog entry: https://micewww.pp.rl.ac.uk/elog/Shift+summary/429
Analysis¶
Some runs seem to have missing spills (10022 - 10038). It may be because ISIS beam had periods of instabilities during these runs. Also, these runs had to be taken with target depth of around 34/35 mm.
Updated by Mohayai, Tanaz Angelina over 5 years ago · 19 revisions